OpenStudy (anonymous):

Simplify -2y8(y3 + 3y4 - 4)

OpenStudy (anonymous):

3y - 4/5 = 1/3y Subtract 1/3y from both sides 8/3y - 4/5 = 0 Add 4/5 to both sides 8/3y = 4/5 Simplify y = 3/10 ANSWER

OpenStudy (anonymous):

a.-2y24 - 6y32 - 4 b.-2y24 - 6y32 + 8y8 c.-2y11 - 6y12 - 4 d.-2y11 - 6y12 + 8y8

OpenStudy (anonymous):

OK Wait sorry :)

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

answer=d

OpenStudy (kuoministers):

To simplify it means to expand and u had to give the answer without explaining it :(

OpenStudy (kuoministers):

You would need to multiply-2y8 to y3 and 3y4 and -4 Which would make answer D